# Equilibrium under Time-Inconsistency: A New Existence Theory by Vanishing Entropy Regularization

### 11 Mar 2026 v1

@@ 4 Existence of Equilibrium by Vanishing Entropy Regularization
where the third equality follows from Part (i), (4.7) and $p^{x}\in L^{q}(D_{N})$, the forth equality follows from (4.2) and $|\lambda_{n_{k}}\mathcal{H}(\pi^{n_{k}})|\to 0$, and the fifth inequality follows from (4.7) again. By virtue of (4.9) and (4.10), the result above implies that
@@
for an arbitrary $\varepsilon$. Consequently, $v^{\infty}(t,x)=V^{\pi^{\infty}}(t,x)$ holds as desired.
Step 3. Finally, we wrap up the proof by showing
$\lim_{k\to\infty}\|v^{n_{k}}(\cdot,x)-v^{\infty}(\cdot,x)\|_{{\mathcal{C}}^{1,\alpha}([0,N])}=0\quad\forall N\in\mathbb{N}.$ (4.11)
Because $v^{\infty}=V^{\pi^{\infty}}$ and $v^{n_{k}}=V^{\pi^{n_{k}}}$ with $\pi^{n_{k}},\pi^{\infty}(x)$ independent of $t$. Following the same arguments for (3.22)—(3.24), we conclude that (4.5) holds. Now take $N\in\mathbb{N}$, and pick an arbitrary $x_{0}\in\mathbb{R}^{d}$, by Part (i), $\|\partial_{t}v^{n_{k}}(\cdot,x_{0})\|_{{\mathcal{C}}^{\alpha/2}([0,N])}\leq A^{*}$. Then, for any subsequence $v^{n_{k_{j}}}$ of $v^{n_{k_{j}}}$, there exists a further subsequence, still denoted by $v^{n_{k_{j}}}$, such that $\|\partial_{t}v^{n_{k_{j}}}(t,x_{0})-u(t)\|_{{\mathcal{C}}_{\alpha}([0,N])}\to 0$ for some $u\in{\mathcal{C}}^{1}_{\alpha/2}[0,N]$. As a result, it holds that
$\int_{0}^{a}\partial_{t}v^{\infty}(t,x_{0})dt=v^{\infty}(a,x_{0})-v^{\infty}(0,x_{0})=\lim_{j\to\infty}(v^{n_{k_{j}}}(a,x_{0})-v^{n_{k_{j}}}(0,x_{0}))$
$=\lim_{j\to\infty}\int_{0}^{a}\partial_{t}v^{n_{k_{j}}}(t,x_{0})dt=\int_{0}^{a}u(t)dt\quad\forall a\in[0,N],$
which implies that $\partial_{t}v^{\infty}(t,x_{0})=u(t)$ on $[0,N]$. Consequently, $\|v^{n_{k_{j}}}(\cdot,x_{0})-v^{\infty}(\cdot,x_{0})\|_{{\mathcal{C}}^{1}_{\alpha}([0,N])}\to 0$. As $v^{n_{k_{j}}}$ is a further subsequence of an arbitrarily picked subsequence of $v^{n_{k}}$, we conclude that
$\|v^{n_{k}}(\cdot,x_{0})-v^{\infty}(\cdot,x_{0})\|_{{\mathcal{C}}^{1}_{\alpha}([0,N])}\to 0,$
@@
We finally verify in the next main result that the limiting $\pi^{\infty}$ is an equilibrium in the original problem.
Theorem 4.1.
Let $(v^{\infty},\pi^{\infty})$ be the limit of $(v^{n_{k}},\pi^{n_{k}})_{k\in\mathbb{N}}$ in Lemma 4.1, then $\pi^{\infty}$ is an equilibrium in Definition 2.1 in the original time-inconsistent control problem.
Proof.
We start with some prior estimates of several terms involving $(v^{n_{k}},\pi^{n_{k}})_{k\in\mathbb{N}}$ and $(v^{\infty},\pi^{\infty})$ near $t=0$. For each $n\in\mathbb{N}\cup\{\infty\}$, define
@@
First, note the fact that for any equi-continuous function sequence $g^{n}:U\mapsto\mathbb{R}$ that converges to a limit $g^{\infty}(a)$, it holds that
@@
where the first equality follows from Lemma 4.1 parts (i) and (ii) and $\lim_{k\to\infty}|\lambda_{n_{k}}\mathcal{H}(\pi^{n_{k}})|_{L^{\infty}}(\mathbb{R}^{d})=0$. As a result, there exists a zero measure set $L\subset\mathbb{R}^{d}$ such that
@@
$b^{\pi^{\infty}}(x)D_{x}v^{\infty}(0,x)+r^{\pi^{\infty}}(0,x)=\sup_{\varpi\in{\mathcal{P}}(U)}\left\{\int_{U}\left[b(x,a)D_{x}u(0,x)+r(0,x,a)\right]\varpi(da)\right\},\quad\forall x\in\mathbb{R}^{d}\setminus L.$ (4.12)
Define ${\mathcal{A}}:=\{\varpi:\mathbb{R}^{d}\mapsto{\mathcal{P}}\text{ is measurable}\}$. There exists a finite constant $C>0$ such that
$\sup_{\varpi\in{\mathcal{A}}}b^{\varpi}(x)D_{x}v^{\infty}(t,x)+r^{\varpi}(t,x)\leq\sup_{\varpi\in{\mathcal{A}}}b^{\varpi}(x)D_{x}v^{\infty}(0,x)+r^{\varpi}(0,x)+Ct^{\alpha/2}$ (4.13)
$\leq b^{\pi^{\infty}}(x)D_{x}v^{\infty}(0,x)+r^{\pi^{\infty}}(0,x)+Ct^{\alpha/2}$
$\leq b^{\pi^{\infty}}(x)D_{x}v^{\infty}(t,x)+r^{\pi^{\infty}}(t,x)+2Ct^{\alpha/2},\quad\forall x\in\mathbb{R}^{d}\setminus L,$
where the first and third inequalities follow from (4.5) in Lemma 4.1 part (iii) and the holder continuity of $r$ on $t$, the second inequality follows from (4.12). Using the same argument and taking the same constant $C$ without loss of generality, we deduce that
$\sup_{k\in\mathbb{N}}b^{\pi^{n_{k}}}(x)D_{x}v^{n_{k}}(t,x)+r^{\pi^{n_{k}}}(t,x)+\lambda_{n_{k}}\delta(t)\mathcal{H}(\pi^{n_{k}}(x))$ (4.14)
$\leq\sup_{k\in\mathbb{N}}b^{\pi^{n_{k}}}(x)D_{x}v^{n_{k}}(0,x)+r^{\pi^{n_{k}}}(0,x)+\lambda_{n_{k}}\mathcal{H}(\pi^{n_{k}}(x))+Ct^{\alpha/2},\quad\forall x\in\mathbb{R}^{d}.$
We show that (2.5) holds for $(v^{\infty},\pi^{\infty})$ by contradiction. Suppose that there exist some $A>0$ and $O\subset\mathbb{R}^{d}$ with $\operatorname{\texttt{Leb}}(O)=B>0$ such that
@@
$A\leq$ $\partial_{t}u(0,x)+\frac{1}{2}\operatorname{tr}\left(\sigma\sigma^{T}(x)D^{2}_{x}u(0,x)\right)\sup_{\varpi\in{\mathcal{P}}(U)}\left\{\int_{U}\left[b(x,a)D_{x}u(0,x)+r(0,x,a)\right]\varpi(da)\right\},\quad\forall x\in O.$
We can get uniform Hölder continuity of the second order term. Using the uniform Hölder continuity in $t$ of the PDE for $v^{n_{k}}$ and $v^{\infty}$, we get that the difference is strictly larger than $A/2$ on a domain $0,\varepsilon]\times B_{r}(x)$, which contradicts with the convergence in $W^{1,2}_{p}$.
@@
Fix an arbitrary $x\in{\mathbb{R}^{d}}$ and take an arbitrary $\varpi\in{\mathcal{A}}$. By the fact $b^{\varpi}\in L^{\infty}({\mathbb{R}^{d}})$ and conditions on $\sigma$ stated in Assumption 2.2, the SDE
$dX^{\varpi}_{s}=b^{\varpi}(X_{s}^{\varpi})dt+\sigma(X^{\varpi}_{s})dW_{s}\quad\text{with }X^{\varpi}_{0}=x$
@@
admits a unique strong solution. Moreover, $X^{\varpi}_{s}$ has a locally Hölder continuous density for any $s>0$. Fix an arbitrary $0<\varepsilon_{0}<1$ and let $\rho_{N}:=\inf_{s\geq 0}\{X^{\varpi}_{s}\notin B_{N}(x)\}$. By (2.1), for any $\epsilon$, there exists $N_{0}>0$ such that
$\mathbb{P}_{x}(\rho_{N}\leq\varepsilon_{0})\leq\frac{\epsilon}{A^{*}+\int_{0}^{\varepsilon_{0}}\sup_{x\in\mathbb{R}^{d},a\in U}|r(s,x,a)|ds},$
for all $N\geq N_{0}$. Noting that $\|v^{\infty}\|_{L^{\infty}(\mathbb{T}\times\mathbb{R}^{d})}\leq A^{*}$, we get that
$\left|\mathbb{E}_{x}[v^{\infty}({\varepsilon_{0}},X^{\varpi}_{{\varepsilon_{0}}})]-\mathbb{E}_{x}[v^{\infty}({\varepsilon_{0}}\wedge\rho_{N},X^{\varpi}_{{\varepsilon_{0}}\wedge\rho_{N}})]\right|$ (4.15)
@@
$\leq\left|\mathbb{E}_{x}\left[v^{\infty}({\varepsilon_{0}},X^{\varpi}_{\varepsilon_{0}}){\mathbf{1}}_{\{\rho_{N}>{\varepsilon_{0}}\}}\right]\right|+\left|\mathbb{E}_{x}\left[\int_{0}^{\varepsilon_{0}}r^{\varpi}(s,X^{\varpi}_{s}){\mathbf{1}}_{\{\rho_{N}>{\varepsilon_{0}}\}}ds\right]\right|$
@@
$\leq\left(\|v^{\infty}\|_{L^{\infty}(\mathbb{T}\times\mathbb{R}^{d})}+\int_{0}^{\varepsilon_{0}}\sup_{x\in\mathbb{R}^{d},a\in U}|r(s,x,a)|ds\right)\mathbb{P}_{x}(\rho_{N}>{\varepsilon_{0}})\leq\epsilon,\quad\forall N\geq N_{0}.$
@@
Take any $N\in\mathbb{N}$ with $N\geq N_{0}$. In view that $v^{\infty}\in W^{1,2}_{p}(D_{N})\cap{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(D_{N})$, applying the generalized Itô-Krylov formula to $v^{\infty}(s,X^{\varpi}_{s})$ (see e.g. [16, Section 2.10, Theorem 1] in a more general setting) yields
$dv^{\infty}(s,X^{\varpi}_{s})$
@@
$=\left(\partial_{t}v^{\infty}(s,X^{\varpi}_{s})+b^{\varpi}(x)D_{x}v^{\infty}(s,X^{\varpi}_{s})+\frac{1}{2}\operatorname{tr}\bigg((\sigma\sigma^{T})(X^{\varpi}_{s})D^{2}_{x}v^{\infty}(s,X^{\varpi}_{s})\bigg)\right)ds$
$\qquad+\sigma(X^{\varpi}_{s})D_{x}v^{\infty}(s,X^{\varpi}_{s})dW_{s}.$
Denote by $p^{x}(s,y):=\mathbb{P}(X^{\varpi}_{s}\in dy|s\leq\rho_{N})$, which belongs to $L^{q}(D_{N})$. Taking expectations on both sides of above equality yields
$\mathbb{E}_{x}[V^{\varpi\otimes_{\varepsilon_{0}}\pi^{\infty}}(\varepsilon\wedge\rho_{N},X^{\varpi}_{\varepsilon_{0}\wedge N})]-J^{\pi^{\infty}}(x)+\mathbb{E}_{x}\left[\int_{0}^{{\varepsilon_{0}}\wedge\rho_{N}}r^{\varpi}(s,X^{\varpi}_{s})ds\right]$ (4.16)
@@
$=\mathbb{E}_{x}\bigg[\int_{0}^{\varepsilon_{0}\wedge\rho_{N}}\bigg(v^{\infty}_{t}(s,X^{\varpi}_{s})+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(X^{\varpi}_{s})D^{2}_{x}v^{\infty}(s,X^{\varpi}_{s})\right)$
$\qquad\qquad\qquad\quad+b^{\varpi}(X^{\varpi}_{s})D_{x}v^{\infty}(s,X^{\varpi}_{s})+r^{\varpi}(s,X^{\varpi}_{s})\bigg)ds\bigg]$

### 04 Apr 2026 v2

@@ 4 Existence of Equilibrium by Vanishing Entropy Regularization
where the third equality follows from Part (i) and the convergence of $b^{\pi^{n_{k}}}$ to $b^{\pi^{\infty}}$ in (4.6), the forth equality follows from (4.4) and $\|\lambda_{n_{k}}\mathcal{H}(\pi^{n_{k}})\|_{L^{\infty}(\mathbb{R}^{d})}\to 0$, and the fifth inequality follows from the convergence of $r^{\pi^{n_{k}}}$ to $r^{\pi^{\infty}}$ in (4.6). By virtue of (4.9) and (4.10), the result above implies that
@@
for an arbitrary $\varepsilon$. Consequently, $v^{\infty}(t,x)=V^{\pi^{\infty}}(t,x)$ holds as desired. Moreover, following the same arguments for (3.28)–(3.30), we conclude that
$\sup_{x\in\mathbb{R}^{d}}\|v^{\infty}(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T})}\leq A^{*},$
@@
Proof of Theorem 4.2:.
We now show that $\pi^{\infty}$ in Lemma 4.1 satisfies the statement in Theorem 4.2. It is directly from Lemma 4.1 that $v^{\infty}=V^{\pi^{\infty}}\in{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})\cap W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})$ with $\sup_{x\in\mathbb{R}^{d}}\|v^{\infty}(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T}\times\mathbb{R}^{d})}<\infty$. Hence, it suffices to show that $v^{\infty}$ is a strong solution to the system:
$+\sup_{\varpi\in{\mathcal{P}}(U)}\left\{\int_{U}\left[b(x,a)D_{x}u(0,x)+r(0,x,a)\right]\varpi(da)\right\},\;\text{a.e on $\mathbb{R}^{d}$},$ (4.11)
$0=$ $\partial_{t}u(t,x)+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(x)D^{2}_{x}u(t,x)\right)+b^{\pi^{\infty}}(x)D_{x}u(t,x)+r^{\pi^{\infty}}(t,x)\;\text{a.e. on $\mathbb{T}\times\mathbb{R}^{d}$}.$ (4.12)
We first show that $v^{\infty}$ is a strong solution to (4.12) by taking limit in (4.4). Notice that $v^{\infty}\in{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})\cap W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})$ and $\sup_{x\in\mathbb{R}^{d}}\|v^{\infty}(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T})}\leq A^{*}$. Fix an arbitrary $(t_{0},x_{0})\in\mathbb{T}\times\mathbb{R}^{d}$. A direct calculation gives that
$0=$ $\lim_{k\to\infty}\int_{D_{1}(t_{0},x_{0})}\bigg(\partial_{t}v^{n_{k}}(t,x)+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(x)D^{2}_{x}v^{n_{k}}(t,x)\right)$
$\qquad\qquad\qquad\qquad+b^{\pi^{n_{k}}}(x)D_{x}v^{n_{k}}(t,x)+r^{\pi^{n_{k}}}(t,x)+\lambda\delta(t)\mathcal{H}(\pi^{n_{k}}(x))\bigg)\phi(t,x)dtdx$
$=$ $\int_{D_{1}(t_{0},x_{0})}\bigg(\partial_{t}v^{\infty}(t,x)+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(x)D^{2}_{x}v^{\infty}(t,x)\right)$
$\qquad\qquad\qquad+b^{\pi^{\infty}}(x)D_{x}v^{\infty}(t,x)+r^{\pi^{\infty}}(t,x)\bigg)\phi(t,x)dtdx,\quad\forall\phi\in L^{q}(D_{1}(t_{0},x_{0})),$
where the second equality follows from $\lim_{k\to\infty}\|\lambda_{n_{k}}\mathcal{H}(\pi^{n_{k}})\|_{L^{\infty}(\mathbb{R}^{d})}=0$, Lemma 4.1 Part (i) and (4.6) in Lemma 4.1 Part (ii). Then by the arbitrariness of $(t_{0},x_{0})$, we conclude that $v^{\infty}$ is a strong solution to (4.12).
Next we show $v^{\infty}$ satisfies (4.11) by contradiction. We start with some a priori estimates involving $(v^{n_{k}},\pi^{n_{k}})_{k\in\mathbb{N}}$ and $(v^{\infty},\pi^{\infty})$ near $t=0$. For each $n\in\mathbb{N}\cup\{\infty\}$, define
@@
First, note the fact that for any equi-continuous function sequence $g^{n}:U\to\mathbb{R}$ that converges to a limit $g^{\infty}(a)$, it holds that
@@
where the first equality follows from Lemma 4.1 Parts (i) and (ii) and $\lim_{k\to\infty}\|\lambda_{n_{k}}\mathcal{H}(\pi^{n_{k}})\|_{L^{\infty}}(\mathbb{R}^{d})=0$. As a result, there exists a zero measure set $L_{0}\subset\mathbb{R}^{d}$ such that
$b^{\pi^{\infty}}(x)D_{x}v^{\infty}(0,x)+r^{\pi^{\infty}}(0,x)$ (4.13)
@@
$=\sup_{\varpi\in{\mathcal{P}}(U)}\left\{\int_{U}\left[b(x,a)D_{x}v^{\infty}(0,x)+r(0,x,a)\right]\varpi(da)\right\},\quad\forall x\in\mathbb{R}^{d}\setminus L_{0}.$
Note that $\pi^{\infty}$ is independent of $t$. Then by
$\|v^{\infty}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\vee\sup_{x\in\mathbb{R}^{d}}\|v^{\infty}(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T}\times\mathbb{R}^{d})}\leq A^{*},$ (4.14)
we conclude that the three terms, $\partial_{t}v^{\infty}(t,x)$, $b^{\pi^{\infty}}(x)D_{x}v^{\infty}(t,x)$, $r^{\pi^{\infty}}(t,x)$, are $\alpha/2$-Hölder continuous on variable $t$ for each $x$. This together with (4.12) yields that the term $\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(x)D^{2}_{x}v^{\infty}(t,x)\right)$ shares the same $\alpha/2$-Hölder continuity on variable $t$ for a.e. $x\in\mathbb{R}^{d}$. Define
$H(t,x):=\partial_{t}v^{\infty}(t,x)+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(x)D^{2}_{x}v^{\infty}(t,x)\right)+b^{\pi^{\infty}}(x)D_{x}v^{\infty}(t,x)+r^{\pi^{\infty}}(t,x)$ (4.15)
We conclude from the above analysis that there exists a zero measure set $L\supset L_{0}$ and a constant $C$ such that for all $x\in\mathbb{R}^{d}\setminus L$,
$|H(t,x)-H(0,x)|\leq Ct^{\alpha/2}\quad\forall t\in[0,1].$ (4.16)
(a). Suppose there exists an open set $B_{r}(x_{0})$ and a constant $A>0$ such that, for all $x\in B_{r}(x_{0})$,
@@
$\partial_{t}v^{\infty}(0,x)+\frac{1}{2}\operatorname{tr}\left(\sigma\sigma^{T}(x)D^{2}_{x}v^{\infty}(0,x)\right)+\sup_{\varpi\in{\mathcal{P}}(U)}\left\{\int_{U}\left[b(x,a)D_{x}v^{\infty}(0,x)+r(0,x,a)\right]\varpi(da)\right\}\geq A.$
Then, we can conclude from (4.13) and (4.16) that there exists $0<\varepsilon_{0}<1$ with
$H(t,x)\geq H(0,x)-\frac{A}{2}\geq A-\frac{A}{2}=\frac{A}{2}\quad\forall(t,x)\in[0,\varepsilon_{0}]\times(B_{r}(x_{0})\setminus L),$
which contradicts with (4.12).
(b). On the other-hand side, if there exists an open set $B_{r}(x_{0})$ and a constant $A>0$ such that, for all $x\in B_{r}(x_{0})$,
$\partial_{t}v^{\infty}(0,x)+\frac{1}{2}\operatorname{tr}\left(\sigma\sigma^{T}(x)D^{2}_{x}v^{\infty}(0,x)\right)+\sup_{\varpi\in{\mathcal{P}}(U)}\left\{\int_{U}\left[b(x,a)D_{x}v^{\infty}(0,x)+r(0,x,a)\right]\varpi(da)\right\}\leq-A.$
Then, we can conclude from (4.16) that there exists $0<\varepsilon<1$ with
$H(t,x)\leq H(0,x)+\frac{A}{2}\leq-A+\frac{A}{2}=-\frac{A}{2}\quad\forall(t,x)\in[0,\varepsilon_{0}]\times(B_{r}(x_{0})\setminus L),$
which again contradicts with (4.12). Then cases (a) and (b) together show that (4.11) holds, which completes the proof. ∎
4.2 Proof of Theorem 4.1
We finally prove the sufficient condition for the existence of equilibrium in Theorem 4.1. As a result, the limiting feedback strategy $\pi^{\infty}$ is an equilibrium that satisfies Definition 2.1.
Proof of Theorem 4.1:.
Take a feedback strategy $\pi^{*}:\mathbb{R}\to{\mathcal{P}}(U)$. Suppose $V^{\pi^{*}}\in{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})\cap W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})$ with $\sup_{x\in\mathbb{R}^{d}}\|V^{\pi^{*}}(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T}\times\mathbb{R}^{d})}<\infty$, and $V^{\pi^{*}}$ is a strong solution to the EHJB system (4.1)–(4.2). Then by the same argument between (4.14) and (4.16), we conclude that
The terms $\partial_{t}V^{\pi^{*}}(t,x)$, $D_{x}V^{\pi*}(t,x)$, $\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(x)D^{2}_{x}V^{\pi*}(t,x)\right)$ (4.17)
[…]
and (4.16) holds by replacing $v^{\infty}$ and $\pi^{\infty}$ in (4.15) with $V^{\pi^{*}}$ and $\pi^{*}$ respectively.
We first show
$\sup_{\varpi\in{\mathcal{A}}}b^{\varpi}(x)D_{x}V^{\pi^{*}}(0,x)+r^{\varpi}(0,x)=b^{\pi^{*}}(x)D_{x}V^{\pi*}(0,x)+r^{\pi^{*}}(0,x)\text{ a.e. on $\mathbb{R}^{d}$}.$ (4.18)
Suppose an open set $B_{r}(x_{0})$ and $A>0$ exist such that
$\sup_{\varpi\in{\mathcal{A}}}b^{\varpi}(x)D_{x}V^{\pi^{*}}(0,x)+r^{\varpi}(0,x)-A\geq b^{\pi^{*}}(x)D_{x}V^{\pi*}(0,x)+r^{\pi^{*}}(0,x)\text{ a.e. on $\mathbb{R}^{d}$}.$
Then by (4.17) and $V^{\pi^{*}}$ satisfying (4.1), we conclude that there exists an constant $0<\varepsilon_{0}\leq 1$ such that
$\partial_{t}V^{\pi^{*}}(t,x)+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(x)D^{2}_{x}V^{\pi^{*}}(t,x)\right)+b^{\pi^{*}}(x)D_{x}V^{\pi^{*}}(t,x)+r^{\pi^{*}}(t,x)$
$=H(t)\leq H(0)+\frac{A}{2}\leq-A+\frac{A}{2}=-\frac{A}{2}\quad\text{a.e. on $[0,\varepsilon_{0}]\times B_{r}(x_{0})$,}$
which contradicts with the assumption that $V^{\pi^{*}}$ satisfies (4.2). Thus, (4.18) holds.
Now we show that $\pi^{*}$ satisfies (2.4). Define ${\mathcal{A}}:=\{\varpi:\mathbb{R}^{d}\to{\mathcal{P}}\text{ is measurable}\}$. There exists a finite constant $C>0$ such that for a.e. $x\in\mathbb{R}^{d}$,
$\sup_{\varpi\in{\mathcal{A}}}b^{\varpi}(x)D_{x}V^{\pi^{*}}(t,x)+r^{\varpi}(t,x)\leq\sup_{\varpi\in{\mathcal{A}}}b^{\varpi}(x)D_{x}V^{\pi^{*}}(0,x)+r^{\varpi}(0,x)+Ct^{\alpha/2}$ (4.19)
$=b^{\pi^{*}}(x)D_{x}V^{\pi*}(0,x)+r^{\pi^{*}}(0,x)+Ct^{\alpha/2}$
$\leq b^{\pi^{*}}(x)D_{x}V^{\pi*}(t,x)+r^{\pi^{*}}(t,x)+2Ct^{\alpha/2},\quad\forall 0\leq t\leq 1.$
@@
where the first and third lines follow from the Hölder continuity of $D_{x}V^{\pi^{*}}$ and $r$ with respect to $t$, the second line follows from (4.18). Fix an arbitrary $x\in{\mathbb{R}^{d}}$ and take an arbitrary $\varpi\in{\mathcal{A}}$. By the fact $b^{\varpi}\in L^{\infty}({\mathbb{R}^{d}})$ and conditions on $\sigma$ stated in Assumption 2.2, the SDE
$dX^{\varpi}_{s}=b^{\varpi}(X_{s}^{\varpi})dt+\sigma(X^{\varpi}_{s})dW_{s}\quad\text{with }X^{\varpi}_{0}=x$ (4.20)
@@
admits a unique strong solution. Moreover, $X^{\varpi}_{s}$ has a locally Hölder continuous density for any $s>0$. Fix an arbitrary $0<\varepsilon_{0}<1$ and let $\rho_{N}:=\inf_{s\geq 0}\{X^{\varpi}_{s}\notin B_{N}(x)\}$. By Assumption 2.2, for any $\epsilon$, there exists $N_{0}>0$ such that
$\mathbb{P}_{x}(\rho_{N}\leq\varepsilon_{0})\leq\frac{\epsilon}{2K_{2}}\quad\forall N\geq N_{0}.$
Noting again that $\|V^{\pi*}\|_{L^{\infty}(\mathbb{T}\times\mathbb{R}^{d})}+\int_{0}^{\infty}\sup_{x\in\mathbb{R}^{d},a\in U}|r(s,x,a)|ds\leq 2K_{2}$, we get that
$\left|\mathbb{E}_{x}[V^{\pi*}({\varepsilon_{0}},X^{\varpi}_{{\varepsilon_{0}}})]-\mathbb{E}_{x}[V^{\pi*}({\varepsilon_{0}}\wedge\rho_{N},X^{\varpi}_{{\varepsilon_{0}}\wedge\rho_{N}})]\right|$ (4.21)
@@
$\leq\left|\mathbb{E}_{x}\left[V^{\pi*}({\varepsilon_{0}},X^{\varpi}_{\varepsilon_{0}}){\mathbf{1}}_{\{\rho_{N}>{\varepsilon_{0}}\}}\right]\right|+\left|\mathbb{E}_{x}\left[\int_{0}^{\varepsilon_{0}}r^{\varpi}(s,X^{\varpi}_{s}){\mathbf{1}}_{\{\rho_{N}>{\varepsilon_{0}}\}}ds\right]\right|$
@@
$\leq\left(\|V^{\pi*}\|_{L^{\infty}(\mathbb{T}\times\mathbb{R}^{d})}+\int_{0}^{\varepsilon_{0}}\sup_{x\in\mathbb{R}^{d},a\in U}|r(s,x,a)|ds\right)\mathbb{P}_{x}(\rho_{N}>{\varepsilon_{0}})\leq\epsilon\quad\forall N\geq N_{0}.$
@@
Take any $N\in\mathbb{N}$ with $N\geq N_{0}$. In view that $V^{\pi*}\in W^{1,2}_{p}(D_{N})\cap{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(D_{N})$, applying again the generalized Itô-Krylov formula to $V^{\pi*}(s,X^{\varpi}_{s})$ yields
$dV^{\pi^{*}}(s,X^{\varpi}_{s})$
@@
$=\left(\partial_{t}V^{\pi*}(s,X^{\varpi}_{s})+b^{\varpi}(x)D_{x}V^{\pi*}(s,X^{\varpi}_{s})+\frac{1}{2}\operatorname{tr}\bigg((\sigma\sigma^{T})(X^{\varpi}_{s})D^{2}_{x}V^{\pi*}(s,X^{\varpi}_{s})\bigg)\right)ds$
$\qquad+\sigma(X^{\varpi}_{s})D_{x}V^{\pi*}(s,X^{\varpi}_{s})dW_{s}.$
Taking expectations on both sides of above equality yields
$\mathbb{E}_{x}[V^{\varpi\otimes_{\varepsilon_{0}}\pi^{*}}(\varepsilon_{0}\wedge\rho_{N},X^{\varpi}_{\varepsilon_{0}\wedge\rho_{N}})]-J^{\pi^{*}}(x)+\mathbb{E}_{x}\left[\int_{0}^{{\varepsilon_{0}}\wedge\rho_{N}}r^{\varpi}(s,X^{\varpi}_{s})ds\right]$ (4.22)
@@
$=\mathbb{E}_{x}\bigg[\int_{0}^{\varepsilon_{0}\wedge\rho_{N}}\bigg(\partial_{t}V^{\pi*}(s,X^{\varpi}_{s})+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(X^{\varpi}_{s})D^{2}_{x}V^{\pi*}(s,X^{\varpi}_{s})\right)$
$\qquad\qquad\qquad\quad+b^{\varpi}(X^{\varpi}_{s})D_{x}V^{\pi*}(s,X^{\varpi}_{s})+r^{\varpi}(s,X^{\varpi}_{s})\bigg)ds\bigg]$